Three years ago, Mark was twice as old as Diane. Four years from now, twice Mark's age will be three less than three times Diane's age. How old are they now?
Please help me solve this problem.
Thank you very much for your cooperation.
Email: kwame730@gmail.com

Answer by ankor@dixie-net.com(22740)

let m = Marks present age
let d = diane's
:
Write an equation for each statement, simplify as much as possible
:
Three years ago, Mark was twice as old as Diane.
m - 3 = 2(d - 3)
m - 3 = 2d - 6
m = 2d - 6 + 3
m = 2d - 3
:
Four years from now, twice Mark's age will be three less than three times Diane's age.
2(m+4) = 3(d+ 4) - 3
2m + 8 = 3d + 12 - 3
2m = 3d + 9 - 8
2m = 3d + 1
replace m with (2d - 3) from the first statement
2(2d-3) = 3d + 1
4d - 6 = 3d + 1
4d - 3d = 6 + 1
d = 7 yrs is Diane's present age
then
m = 2(7) - 3
m = 11 yrs is Marks age
:
;
You can check this for yourself in both statements